# Chow groups of one-dimensional noetherian domains

### v2

@@ 1. Introduction
of all invertible fractional ideals modulo the principal fractional ideals of $\mathcal{O}$. In the case $\mathcal{O}={\tilde{\mathcal{O}}}$ this group is called the class group, and every non-zero fractional ideal of ${\tilde{\mathcal{O}}}$ is invertible and can be uniquely written as a product of maximal ideals. A fundamental result of algebraic number theory states that the group $\operatorname{Pic}({\tilde{\mathcal{O}}})$ is always finite if $K$ is a global field. We denote by the class number the order of the class group. We get that ${\tilde{\mathcal{O}}}$ is a principal ideal domain if and only if it has class number 1. The behavior of class groups is still mysterious and a famous open problem is the so-called class number 1 problem which is the question, if there exist infinitely many number fields with class number 1. On the other hand we know many infinite families of fields which have non-trivial class numbers, e.g. there are only finitely many imaginary quadratic number fields which have class number 1.
Assuming that we have computed the unit and class groups of ${\tilde{\mathcal{O}}}$, [4] gives algorithms to compute the unit and Picard groups of $\mathcal{O}$. The Picard group gives no information about the structure of non-invertible ideals. The non-invertible prime ideals of $\mathcal{O}$ are exactly the (finitely many) prime ideals of $\mathcal{O}$ which contain the conductor :=$\mathcal{F}:=\{x\in K\mid x{\tilde{\mathcal{O}}}\subseteq\mathcal{O}\}$. This is the largest ideal of ${\tilde{\mathcal{O}}}$ that is contained in $\mathcal{O}$. It is very classical (see Definition 2.4) to define the Chow group
@@
The goal of this paper is to understand the behavior of Chow groups and its relation to $\operatorname{Pic}(\mathcal{O})$ or $\operatorname{Pic}({\tilde{\mathcal{O}}})$. We are also interested in the relation between the Chow groups of two orders $\mathcal{O}^{\prime}\subseteq\mathcal{O}$. For orders in a global field $K$, we present an algorithm to compute the Chow group of $\mathcal{O}$.
@@
$0\rightarrow\operatorname{Pic}({\tilde{\mathcal{O}}})/R\rightarrow\operatorname{Chow}(\mathcal{O})\rightarrow\bigoplus_{i=1}^{r}\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}_{i}})\rightarrow 0,$
where $\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}_{i}})$ is the Chow group of the localization of $\mathcal{O}$ at $\mathfrak{p}_{i}$. The local Chow groups are cyclic groups and the order can be computed by considering the inertia degree of $\mathfrak{p}_{i}$ and the inertia degrees of the prime ideals of ${\tilde{\mathcal{O}}}$ lying above $\mathfrak{p}_{i}$. In Theorem 2.9 we also show that the group of relations $R$ can be computed by considering all prime ideals of ${\tilde{\mathcal{O}}}$ which lie over one of the prime ideals $\mathfrak{p}_{i}$. We remark that the above sequence is not necessarily split, see Example 4.1.
There is a canonical homomorphism from $\operatorname{Pic}(\mathcal{O})\rightarrow\operatorname{Chow}(\mathcal{O})$ and we give exact criteria when this homomorphism is injective or surjective in Proposition 2.14. Similarly, we discuss the canonical homomorphism $\operatorname{Chow}({\tilde{\mathcal{O}}})\rightarrow\operatorname{Chow}(\mathcal{O})$, which is in general neither injective nor surjective.
@@
The paper is organized as follows. In Section 2, we recall basic definitions regarding divisors and Chow groups. We discuss canonical homomorphisms between such groups belonging to different rings. In Theorem 2.11 we show how to compute Chow groups and in Algorithm 2.12 we give a principal divisor test. We end this section by analysing the homomorphism between the Picard and Chow group of $\mathcal{O}$. Section 3 shows that many number fields contain orders with trivial Chow group and Section 4 gives some examples. The last section corrects some wrong statements from [9]. In Theorem 5.1 we describe all $\mathcal{O}$ such that $\mathcal{O}^{*}={\tilde{\mathcal{O}}}^{*}$ and $\operatorname{Pic}(\mathcal{O})\cong\operatorname{Pic}({\tilde{\mathcal{O}}})$. In Theorem 5.4 we show that the canonical homomorphism $\mathcal{I}(\mathcal{O})\to\operatorname{Div}(\mathcal{O})$ is injective if and only if $\mathcal{O}={\tilde{\mathcal{O}}}$.
@@ Acknowledgments
The authors thank Stefano Marseglia and the anonymous referee for helpful comments. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-TRR 358/1 2023 – 491392403.
@@ 2. Divisors and Chow groups
By [8, Lemma I.12.7], the degree $\deg_{\mathfrak{p}}(\mathfrak{P})$ of $\mathcal{O}^{\prime}/\mathfrak{P}$ over $\mathcal{O}/\mathfrak{p}$ is finite.
For a maximal ideal $\mathfrak{p}\in\mathbb{P}(\mathcal{O})$, let $\mathcal{O}_{\mathfrak{p}}$ be the localization of $\mathcal{O}$ at $\mathfrak{p}$. More generally, if $M$ is an $\mathcal{O}$-module, we denote by $M_{\mathfrak{p}}:=M\otimes_{\mathcal{O}}\mathcal{O}_{\mathfrak{p}}$ the localization of $M$ at $\mathfrak{p}$. We follow Neukirch [8, Section I.12] to define the divisor and Chow groups of $\mathcal{O}$. For $\mathfrak{p}\in\mathbb{P}(\mathcal{O})$ define
@@
Lemma 2.1.
Let $a\in K^{*}$ and let $\mathfrak{P}_{1},\dots,\mathfrak{P}_{r}$ be the prime ideals of $\mathcal{O}^{\prime}$ over $\mathfrak{p}$. Then
Proof.
See [1, Example A.3.1]. ∎
Corollary 2.2.
If $a\in{\tilde{\mathcal{O}}}_{\mathfrak{p}}^{*}$ is a unit in the normalization, then $\operatorname{ord}_{\mathfrak{p}}(a)=0$.
Example 2.3.
Then $a{\tilde{\mathcal{O}}}_{\mathfrak{P}}=\mathfrak{P}{\tilde{\mathcal{O}}}_{\mathfrak{P}}$. Hence $\operatorname{ord}_{\mathfrak{P}}(a)=1$ and similarly $\operatorname{ord}_{\overline{\mathfrak{P}}}(a)=-1$. Lemma 2.1 shows that for $\mathfrak{p}=\mathfrak{P}\cap\mathcal{O}=\overline{\mathfrak{P}}\cap\mathcal{O}$ we find $\operatorname{ord}_{\mathfrak{p}}(a)=1-1=0$.
@@
Usually, the Chow group of a one-dimensional scheme is defined by 00-cycles modulo rational equivalence to 00, see [1, Section 1.3]. The 00-cycles correspond to divisors in $\operatorname{Div}(\mathcal{O})$ and this correspondence identifies principal divisors with 00-cycles which are rationally equivalent to 00.
Remark 2.5.
By [8, Proposition I.12.8], the normalization ${\tilde{\mathcal{O}}}$ is a Dedekind ring. Hence the (multiplicative) group of fractional ${\tilde{\mathcal{O}}}$-ideals $\mathcal{I}({\tilde{\mathcal{O}}})$ is freely generated by the set of maximal ideals $\mathbb{P}({\tilde{\mathcal{O}}})$. Hence $\operatorname{div}_{\tilde{\mathcal{O}}}\colon\mathcal{I}({\tilde{\mathcal{O}}})\to\operatorname{Div}({\tilde{\mathcal{O}}})$ is an isomorphism that maps principal ideals to principal divisors and vice versa. Therefore it induces a canonical isomorphism
$\overline{\operatorname{div}}_{\tilde{\mathcal{O}}}\colon\operatorname{Pic}({\tilde{\mathcal{O}}})\to\operatorname{Chow}({\tilde{\mathcal{O}}})$
between the Picard group and the Chow group of ${\tilde{\mathcal{O}}}$.
Our description of $\operatorname{Chow}(\mathcal{O})$ is based on the following result.
@@
$f_{*}\colon\operatorname{Div}(\mathcal{O}^{\prime})\to\operatorname{Div}(\mathcal{O}),\;\sum_{\mathfrak{P}\in\mathbb{P}(\mathcal{O}^{\prime})}a_{\mathfrak{P}}\mathfrak{P}\mapsto\sum_{\mathfrak{P}\in\mathbb{P}(\mathcal{O}^{\prime})}a_{\mathfrak{P}}\deg_{f(\mathfrak{P})}(\mathfrak{P})\cdot f(\mathfrak{P})$
mapping principal divisors to principal divisors. Hence it induces a homomorphism
@@
See [1, Theorem 1.4] or [10, Lemma 42.18.1]. ∎
We will show in Example 4.2 that the homomorphism $\operatorname{Chow}(\mathcal{O}^{\prime})\to\operatorname{Chow}(\mathcal{O})$ is in general neither injective nor surjective.
Remark 2.7.
Lemma 2.1 shows that for $a\in K^{*}$ we have
(2.1) $f_{*}(\operatorname{div}_{\tilde{\mathcal{O}}}(a))=f_{*}\left(\sum_{\mathfrak{P}\in\mathbb{P}(\mathcal{O}^{\prime})}\operatorname{ord}_{\mathfrak{P}}(a)\mathfrak{P}\right)=\sum_{\mathfrak{p}\in\mathbb{P}(\mathcal{O})}\operatorname{ord}_{\mathfrak{p}}(a)\mathfrak{p}=\operatorname{div}_{\mathcal{O}}(a)\>.$
The conductor :=$\mathcal{F}:=\{x\in K\mid x{\tilde{\mathcal{O}}}\subseteq\mathcal{O}\}$ of $\mathcal{O}$ is the largest ideal of ${\tilde{\mathcal{O}}}$ that is contained in $\mathcal{O}$. A maximal ideal $\mathfrak{p}$ of $\mathcal{O}$ is invertible (or regular) if and only if $\mathfrak{p}$ is coprime to the conductor $\mathcal{F}$. If this is the case, then $\mathfrak{p}{\tilde{\mathcal{O}}}$ is the unique maximal ideal of ${\tilde{\mathcal{O}}}$ over $\mathfrak{p}:=\mathfrak{p}{\tilde{\mathcal{O}}}\cap\mathcal{O}$ and ${\tilde{\mathcal{O}}}_{\mathfrak{p}{\tilde{\mathcal{O}}}}\cong\mathcal{O}_{\mathfrak{p}}$. Thus ${\tilde{\mathcal{O}}}/\mathfrak{p}{\tilde{\mathcal{O}}}\cong\mathcal{O}/\mathfrak{p}$ and therefore $\deg_{\mathfrak{p}}(\mathfrak{p}{\tilde{\mathcal{O}}})=1$. In particular, $f_{*}(\mathfrak{p}\mathcal{O}^{\prime})=\mathfrak{p}$ for any ring $\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq{\tilde{\mathcal{O}}}$. See [8, Proposition I.12.10] for details.
@@
    $\operatorname{im}(\overline{f_{*}})=\operatorname{im}(f_{*})/\mathcal{P}(\mathcal{O})\quad\mbox{and}\quad\ker(\overline{f_{*}})=\{[x]\mid x\in\ker(f_{*})\}\>.$
@@
Part (1) and $\operatorname{im}(\overline{f_{*}})$ follow from Proposition 2.6 and the preceding discussion. If $x\in\ker(f_{*})$ then $[x]\in\ker(\overline{f_{*}})$. Suppose $[x]\in\operatorname{Chow}(\mathcal{O}^{\prime})$ lies in the kernel of $\overline{f_{*}}$. Then there exists some $a\in K^{*}$ such that
[…]
Hence $x-\sum_{\mathfrak{P}\in\mathbb{P}(\mathcal{O}^{\prime})}\operatorname{ord}_{\mathfrak{P}}(a)\mathfrak{P}$ represents $[x]$ and lies in the kernel of $f_{*}$. To prove the last assertion, consider the commutative diagram
@@
with integers $\lambda_{i,j}$. Then Corollary 2.8 yields the exact sequence
$0\rightarrow\operatorname{im}(\overline{f_{*}})\rightarrow\operatorname{Chow}(\mathcal{O})\rightarrow\operatorname{coker}{\overline{f_{*}}}\cong\bigoplus_{i=1}^{r}\mathbb{Z}/g_{i}\mathbb{Z}\rightarrow 0\>.$
Since ${\tilde{\mathcal{O}}}_{\mathfrak{p}}$ has only finitely many prime ideals, it is a principal ideal domain. The preceding discussion yields the following result.
@@
    We have an exact sequence
    $0\rightarrow\operatorname{im}(\overline{f_{*}})\rightarrow\operatorname{Chow}(\mathcal{O})\rightarrow\bigoplus_{i=1}^{r}\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}_{i}})\rightarrow 0\>,$
    where the third homomorphism is the canonical projection
    $[\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i}+\sum_{\mathfrak{q}}a_{\mathfrak{q}}\mathfrak{q}]\mapsto([a_{1}\mathfrak{p}_{1}],\dots,[a_{r}\mathfrak{p}_{r}]),$
    where $\mathfrak{q}$ runs over the invertible prime ideals of $\mathcal{O}$.
Hence $\operatorname{Chow}(\mathcal{O})$ is an extension of $\operatorname{im}(\overline{f_{*}})\cong\operatorname{Pic}({\tilde{\mathcal{O}}})/\{[\mathfrak{A}]\mid\operatorname{div}_{\tilde{\mathcal{O}}}(\mathfrak{A})\in\ker(f_{*})\}$ by the direct sum of the local Chow groups $\bigoplus_{i=1}^{r}\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}_{i}})\cong\bigoplus_{i=1}^{r}\mathbb{Z}/g_{i}\mathbb{Z}$.
In Example 4.1 we will see that this extension is not split in general.
We close this section with two algorithms. To this end, we start with a remark on how to compute the kernel $\ker(f_{*})$.
Remark 2.10.
For $1\leq i\leq r$ let $\mathfrak{Q}_{i}=\sum_{j=1}^{r_{i}}\lambda_{i,j}\mathfrak{P}_{i,j}$, where the $\lambda_{i,j}$ are defined in Eq. (2.2). Then the kernel of $f_{*}$ is generated by
@@
Proof.
The given generators lie in $\ker(f_{*})$. Conversely, let $D:=\sum_{i,j}a_{i,j}\mathfrak{P}_{i,j}\in\ker(f_{*})$. Then $\sum_{j}a_{i,j}d_{i,j}=0$ for all $i$, see Corollary 2.8. Therefore
$D+\sum_{i,j}a_{i,j}(\tfrac{d_{i,j}}{g_{i}}\mathfrak{Q}_{i}-\mathfrak{P}_{i,j})=\sum_{i,j}\tfrac{a_{i,j}d_{i,j}}{g_{i}}\mathfrak{Q}_{i}=\sum_{i}\tfrac{1}{g_{i}}(\sum_{j}a_{i,j}d_{i,j})\mathfrak{Q}_{i}=0\>.$
The result follows. ∎
Theorem 2.11.
For $1\leq i\leq r$ set $\mathfrak{Q}_{i}=\prod_{j=1}^{r_{i}}\mathfrak{P}_{i,j}^{\lambda_{i,j}}$ and
[…]
Let $G=\bigoplus_{i=1}^{r}\mathbb{Z}\mathfrak{p}_{i}\oplus\operatorname{Pic}({\tilde{\mathcal{O}}})/N$. Then
$\varphi\colon G\to\operatorname{Chow}(\mathcal{O}),\,(\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i},[\mathfrak{A}])\mapsto[\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i}+f_{*}(\operatorname{div}_{\tilde{\mathcal{O}}}(\mathfrak{A}))]$
@@
$R=\langle\{(g_{i}\mathfrak{p}_{i},[\mathfrak{Q}_{i}^{-1}])\in G\mid 1\leq i\leq r\}\rangle\>.$

### v3

@@ 1. Introduction
of all invertible fractional ideals modulo the principal fractional ideals of $\mathcal{O}$. In the case $\mathcal{O}={\tilde{\mathcal{O}}}$ this group is called the class group, and every non-zero fractional ideal of ${\tilde{\mathcal{O}}}$ is invertible and can be uniquely written as a product of maximal ideals. A fundamental result of algebraic number theory states that the group $\operatorname{Pic}({\tilde{\mathcal{O}}})$ is always finite if $K$ is a global field. The class number of $K$ is the order of the class group $\operatorname{Pic}({\tilde{\mathcal{O}}})$. We get that ${\tilde{\mathcal{O}}}$ is a principal ideal domain if and only if it has class number 1. The behavior of class groups is still mysterious and a famous open problem is the so-called class number 1 problem which is the question, if there exist infinitely many number fields with class number 1. On the other hand we know many infinite families of number fields which have non-trivial class numbers, e.g. there are only finitely many imaginary quadratic number fields which have class number 1.
The Picard group gives no information about the structure of non-invertible ideals. The non-invertible prime ideals of $\mathcal{O}$ are exactly the (finitely many) prime ideals of $\mathcal{O}$ which contain the conductor :=$\mathcal{F}:=\{x\in K\mid x{\tilde{\mathcal{O}}}\subseteq\mathcal{O}\}$. This is the largest ideal of ${\tilde{\mathcal{O}}}$ that is contained in $\mathcal{O}$. It is very classical (see Definition 2.1) to define the Chow group
@@
The goal of this paper is to understand the behavior of Chow groups and its relation to $\operatorname{Pic}(\mathcal{O})$ or $\operatorname{Pic}({\tilde{\mathcal{O}}})$. For example, there is a canonical homomorphism from $\operatorname{Pic}(\mathcal{O})\rightarrow\operatorname{Chow}(\mathcal{O})$ and Proposition 2.10 discusses when this homomorphism is injective or surjective. We are also interested in the relation between the Chow groups of two orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$. There is a canonical homomorphism $\operatorname{Chow}(\mathcal{O}^{\prime})\rightarrow\operatorname{Chow}(\mathcal{O})$ and its image and kernel are given by Lemma 2.5.
@@
$0\rightarrow\operatorname{Chow}({\tilde{\mathcal{O}}})/R\rightarrow\operatorname{Chow}(\mathcal{O})\rightarrow\bigoplus_{i=1}^{r}\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}_{i}})\rightarrow 0,$
where $\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}_{i}})$ is the Chow group of the localization of $\mathcal{O}$ at $\mathfrak{p}_{i}$ and $R$ denotes the kernel of the canonical homomorphism $\operatorname{Chow}({\tilde{\mathcal{O}}})\to\operatorname{Chow}(\mathcal{O})$. The local Chow groups are cyclic groups and the order can be computed by considering the degree of the residue field $\mathcal{O}/\mathfrak{p}_{i}$ and the degrees of the residue fields of prime ideals of ${\tilde{\mathcal{O}}}$ lying above $\mathfrak{p}_{i}$. In Theorem 2.6 we also show that the group of relations $R$ can be computed by considering all prime ideals of ${\tilde{\mathcal{O}}}$ which lie over one of the prime ideals $\mathfrak{p}_{i}$. We remark that the above sequence is not necessarily split, see Example 4.1.
In the number field case, [5] gives an algorithm to compute $\operatorname{Pic}(\mathcal{O})$ from $\operatorname{Pic}({\tilde{\mathcal{O}}})$. In the same vein, we show in Section 2 how to compute $\operatorname{Chow}(\mathcal{O})$ from $\operatorname{Chow}({\tilde{\mathcal{O}}})\cong\operatorname{Pic}({\tilde{\mathcal{O}}})$ provided we can factor the conductor of $\mathcal{O}$. It is well known how to do this in global fields.
Note that for a global function field $K/\mathbb{F}_{q}(t)$, we can consider the Chow group $\operatorname{Chow}(K)$ of the corresponding projective smooth curve. Following [10, ch. 14] denote by $S_{K}$ all primes of $K$ and by $S$ the finitely many primes of $K$ above the infinite prime of $\mathbb{F}_{q}(t)$. Then the $S$-order $\mathcal{O}_{S}:=\{a\in K\mid\operatorname{ord}_{\mathfrak{P}}(a)\geq 0,\;\forall\mathfrak{P}\in S_{K}\setminus S\}$ is our maximal order $\tilde{\mathcal{O}}$. Then $\operatorname{Chow}(K)$ is isomorphic to $\mathbb{Z}+\operatorname{Chow}^{0}(K)$, where $\operatorname{Chow}^{0}(K)=\operatorname{Div}^{0}(K)/\mathcal{P}(K)$ are the divisors of degree 0 modulo the principal divisors. From [10, Prop. 14.1] we get the exact sequence
$0\rightarrow\operatorname{Div}^{0}(S)/\mathcal{P}(S)\rightarrow\operatorname{Chow}^{0}(K)\rightarrow\operatorname{Chow}(\tilde{\mathcal{O}})\rightarrow C\rightarrow 0,$
where $C$ is a cyclic group of order $d$, and $d$ is the gcd of the degrees of the divisors in $S$. The group $\operatorname{Div}^{0}(S)/\mathcal{P}(S)$ is generated by the degree 0 divisors generated by the primes in $S$. This exact sequence is used in [4, Kor 1.3] to compute $\operatorname{Chow}(\tilde{\mathcal{O}})$ and the corresponding function SClassGroupExactSequence is available in Magma.
@@
The paper is organized as follows. In Section 2, we recall basic definitions regarding divisors and Chow groups. We discuss canonical homomorphisms between such groups belonging to different rings. In Theorem 2.7 we show how to compute Chow groups and in Algorithm 2.8 we give a principal divisor test. We end this section by analyzing the homomorphism between the Picard and Chow group of $\mathcal{O}$. Section 3 shows that many number fields contain orders with trivial Chow group and Section 4 gives some examples. The last section corrects some wrong statements from [11]. In Theorem 5.1 we describe all $\mathcal{O}$ such that $\mathcal{O}^{*}={\tilde{\mathcal{O}}}^{*}$ and $\operatorname{Pic}(\mathcal{O})\cong\operatorname{Pic}({\tilde{\mathcal{O}}})$. In Theorem 5.4 we show that the canonical homomorphism $\mathcal{I}(\mathcal{O})\to\operatorname{Div}(\mathcal{O})$ is injective if and only if $\mathcal{O}={\tilde{\mathcal{O}}}$.
@@ Acknowledgments
The authors thank Stefano Marseglia for helpful comments. We are also grateful to the reviewer for numerous suggestions and the proof of part (2) of Lemma 3.1.
@@ 2. Divisors and Chow groups
By [9, Lemma I.12.7], the degree $\deg_{\mathfrak{p}}(\mathfrak{P})$ of $\mathcal{O}^{\prime}/\mathfrak{P}$ over $\mathcal{O}/\mathfrak{p}$ is finite.
For a maximal ideal $\mathfrak{p}\in\mathbb{P}(\mathcal{O})$, let $\mathcal{O}_{\mathfrak{p}}$ be the localization of $\mathcal{O}$ at $\mathfrak{p}$. More generally, if $M$ is an $\mathcal{O}$-module, we denote by $M_{\mathfrak{p}}:=M\otimes_{\mathcal{O}}\mathcal{O}_{\mathfrak{p}}$ the localization of $M$ at $\mathfrak{p}$. We follow Neukirch [9, Section I.12] to define the divisor and Chow groups of $\mathcal{O}$. For $\mathfrak{p}\in\mathbb{P}(\mathcal{O})$ define
@@
More generally, the Chow group of a one-dimensional noetherian scheme is defined by 00-cycles modulo rational equivalence to 00, see [1, Section 1.3]. The 00-cycles correspond to divisors in $\operatorname{Div}(\mathcal{O})$ and this correspondence identifies principal divisors with 00-cycles which are rationally equivalent to 00.
is called the Picard group of $\mathcal{O}$. Being invertible is a local property. For an invertible fractional ideal $\mathfrak{a}$ of $\mathcal{O}$ and $\mathfrak{p}\in\mathbb{P}(\mathcal{O})$ there exists some $a_{\mathfrak{p}}\in K^{*}$ such that $\mathfrak{a}_{\mathfrak{p}}=a\mathcal{O}_{\mathfrak{p}}$, see [9, Satz (12.4)] for details. Hence we can define
which induces a homomorphism
The groups $\mathcal{I}({\tilde{\mathcal{O}}})$ and $\operatorname{Div}({\tilde{\mathcal{O}}})$ are freely generated by the maximal ideals of ${\tilde{\mathcal{O}}}$ hence $\operatorname{div}_{\tilde{\mathcal{O}}}\colon\mathcal{I}({\tilde{\mathcal{O}}})\to\operatorname{Div}({\tilde{\mathcal{O}}})$ and $\overline{\operatorname{div}}_{\tilde{\mathcal{O}}}\colon\operatorname{Pic}({\tilde{\mathcal{O}}})\to\operatorname{Chow}({\tilde{\mathcal{O}}})$ are canonical isomorphisms.
We now come to a description of $\operatorname{Chow}(\mathcal{O})$, which is based on the following result.
@@
  1. (1)
    Let $a\in K^{*}$ and $\mathfrak{p}\in\mathbb{P}(\mathcal{O})$. If $\mathfrak{P}_{1},\dots,\mathfrak{P}_{r}$ denote the prime ideals of $\mathcal{O}^{\prime}$ over $\mathfrak{p}$, then
  2. (2)
@@
    $f_{*}\colon\operatorname{Div}(\mathcal{O}^{\prime})\to\operatorname{Div}(\mathcal{O}),\;\sum_{\mathfrak{P}\in\mathbb{P}(\mathcal{O}^{\prime})}a_{\mathfrak{P}}\mathfrak{P}\mapsto\sum_{\mathfrak{P}\in\mathbb{P}(\mathcal{O}^{\prime})}a_{\mathfrak{P}}\deg_{f(\mathfrak{P})}(\mathfrak{P})\cdot f(\mathfrak{P})\>.$
    If $a\in K^{*}$, then
    (2.1) $f_{*}(\operatorname{div}_{\mathcal{O}}^{\prime}(a))=f_{*}\left(\sum_{\mathfrak{P}\in\mathbb{P}(\mathcal{O}^{\prime})}\operatorname{ord}_{\mathfrak{P}}(a)\mathfrak{P}\right)=\sum_{\mathfrak{p}\in\mathbb{P}(\mathcal{O})}\operatorname{ord}_{\mathfrak{p}}(a)\mathfrak{p}=\operatorname{div}_{\mathcal{O}}(a)\>.$
    Hence $f_{*}$ maps principal divisors to principal divisors and therefore induces a homomorphism
@@
See [1, Example A.3.1] for a proof of part (1) and [1, Theorem 1.4] or [12, Lemma 42.18.1] for a proof of (2). ∎
We will show in Example 4.2 that the homomorphism $\operatorname{Chow}(\mathcal{O}^{\prime})\to\operatorname{Chow}(\mathcal{O})$ is in general neither injective nor surjective. Part (1) of Proposition 2.2 immediately gives the following result.
@@
Let $a\in{\tilde{\mathcal{O}}}_{\mathfrak{p}}$ be an element in the normalization. Then $a$ is a unit if and only if $\operatorname{ord}_{\mathfrak{p}}(a)=0$.
Example 2.4.
Then $a{\tilde{\mathcal{O}}}_{\mathfrak{P}}=\mathfrak{P}{\tilde{\mathcal{O}}}_{\mathfrak{P}}$. Hence $\operatorname{ord}_{\mathfrak{P}}(a)=1$ and similarly $\operatorname{ord}_{\overline{\mathfrak{P}}}(a)=-1$. Proposition 2.2 shows that for $\mathfrak{p}=\mathfrak{P}\cap\mathcal{O}=\overline{\mathfrak{P}}\cap\mathcal{O}$ we find $\operatorname{ord}_{\mathfrak{p}}(a)=1-1=0$.
The conductor :=$\mathcal{F}:=\{x\in K\mid x{\tilde{\mathcal{O}}}\subseteq\mathcal{O}\}$ of $\mathcal{O}$ is the largest ideal of ${\tilde{\mathcal{O}}}$ that is contained in $\mathcal{O}$. A maximal ideal $\mathfrak{p}$ of $\mathcal{O}$ is invertible (or regular) if and only if $\mathfrak{p}$ is coprime to the conductor $\mathcal{F}$. If this is the case, then $\mathfrak{p}{\tilde{\mathcal{O}}}$ is the unique maximal ideal of ${\tilde{\mathcal{O}}}$ over $\mathfrak{p}=\mathfrak{p}{\tilde{\mathcal{O}}}\cap\mathcal{O}$ and ${\tilde{\mathcal{O}}}_{\mathfrak{p}{\tilde{\mathcal{O}}}}\cong\mathcal{O}_{\mathfrak{p}}$. Thus ${\tilde{\mathcal{O}}}/\mathfrak{p}{\tilde{\mathcal{O}}}\cong\mathcal{O}/\mathfrak{p}$ and therefore $\deg_{\mathfrak{p}}(\mathfrak{p}{\tilde{\mathcal{O}}})=1$. In particular, $f_{*}(\mathfrak{p}\mathcal{O}^{\prime})=\mathfrak{p}$ for any ring $\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq{\tilde{\mathcal{O}}}$. See [9, Proposition I.12.10] for details.
Lemma 2.5.
@@
    $\operatorname{im}(\overline{f_{*}})=\operatorname{im}(f_{*})/\mathcal{P}(\mathcal{O})\quad\mbox{and}\quad\ker(\overline{f_{*}})=(\ker(f_{*})+\mathcal{P}(\mathcal{O}^{\prime}))/\mathcal{P}(\mathcal{O}^{\prime})\>.$
@@
Part (1) follows from Proposition 2.2 and the preceding discussion. Equation (2.1) shows that $f_{*}\colon\mathcal{P}(\mathcal{O}^{\prime})\to\mathcal{P}(\mathcal{O})$ is surjective. This gives the image and kernel of $\overline{f_{*}}$. To prove the last assertion, consider the commutative diagram
@@
with integers $\lambda_{i,j}$. Then Lemma 2.5 yields the exact sequence
$0\rightarrow\operatorname{im}(\overline{f_{*}})\rightarrow\operatorname{Chow}(\mathcal{O})\rightarrow\operatorname{coker}(\overline{f_{*}})\cong\bigoplus_{i=1}^{r}\mathbb{Z}/g_{i}\mathbb{Z}\rightarrow 0\>.$
Since ${\tilde{\mathcal{O}}}_{\mathfrak{p}}$ has only finitely many prime ideals, it is a principal ideal domain. Thus every ideal or divisor of ${\tilde{\mathcal{O}}}_{\mathfrak{p}}$ is principal and thus $\operatorname{Chow}({\tilde{\mathcal{O}}}_{\mathfrak{p}})\to\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}})$ is trivial. Hence the preceding discussion yields the following result.
@@
    The sequence
    $0\rightarrow\operatorname{im}(\overline{f_{*}})\rightarrow\operatorname{Chow}(\mathcal{O})\rightarrow\bigoplus_{\mathfrak{p}\in\mathbb{P}(\mathcal{O})}\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}})\rightarrow 0$
    is exact.
Hence, $\operatorname{Chow}(\mathcal{O})$ is an extension of $\operatorname{im}(\overline{f_{*}})\cong\operatorname{Chow}({\tilde{\mathcal{O}}})/\ker(\overline{f_{*}})$ by the direct sum of the local Chow groups $\bigoplus_{i=1}^{r}\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}_{i}})\cong\bigoplus_{i=1}^{r}\mathbb{Z}/g_{i}\mathbb{Z}$.
The following theorem describes this extension, which is not necessarily split, c.f. Example 4.1.
Theorem 2.7.
For $1\leq i\leq r$ let $Q_{i}=\sum_{j=1}^{r_{i}}\lambda_{i,j}\mathfrak{P}_{i,j}\in\operatorname{Div}({\tilde{\mathcal{O}}})$, where the $\lambda_{i,j}$ are defined in Equation (2.2).
  1. (1)
    The kernel of $f_{*}$ is generated by
@@
  2. (2)
    Let $G=\bigoplus_{i=1}^{r}\mathbb{Z}\mathfrak{p}_{i}\oplus\operatorname{Chow}({\tilde{\mathcal{O}}})/\ker(\overline{f_{*}})$. Then
    $\varphi\colon G\to\operatorname{Chow}(\mathcal{O}),\,(\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i},[A])\mapsto[\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i}+f_{*}(A))]$
@@
    $R=\langle\{(g_{i}\mathfrak{p}_{i},[-Q_{i}])\in G\mid 1\leq i\leq r\}\rangle\>.$